puppylover123:

If a + 1 = b, then b > a. A. Create a pair of values for a and b in which the numerical relationship shown in the given conditional statement is true. In your final answer, include the values of the variables and proof of truth. B. Is it possible to create a pair of values for a and b in which the numerical relationship shown in the given conditional statement is false? Use complete sentences to explain your answer.

jhonyy9:

so for point A. just give values to ,,a" and ,,b" and look what will get for example a = 1 1 + 1 = 2 => 2 > 1 so b > a a = 2 2 + 1 = 3 => 3 > 2 hope these help you @XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so for point A. just give values to ,,a" and ,,b" and look what will get for example a = 1 1 + 1 = 2 => 2 > 1 so b > a a = 2 2 + 1 = 3 => 3 > 2 hope these help you @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) Great job!